Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Solve with combinations

Expert replies
by Jinglander » Tue Sep 28, 2010 8:32 am
I know how to solve this using a probability decision tree, but wanted to see if it could be solved with combinations.

Renee, has a bag of 6 candies, 4 are sweet 2 are sour. Jack picks two candies simultaneously and at random, what is the chance that exactly one candy he picked is sour.

I know the total combinations its just 6!/2!4!. But how do I find the number of outcomes that have exactly one sour. Clearly the numbers are small enough that I can just count the number of out comes but i want a formula that can be applied to larger numbers
Join the discussion
Source: — Problem Solving |

by Geva@EconomistGMAT » Tue Sep 28, 2010 9:24 am
Caveat: the method below is a rather long and complicated approach, but that's what you asked for. What I would do with your original problem is find probability one step at a time. Split into separate events: first choice, second choice.
Probability of exactly one sour means two scenarios: sour only on first, or sour only on second.
P(sour on first choice) 2/6 * P(not sour on second choice) 4/5 = 8/30.
OR
P(not sour on first choice) 4/6 * P(sour on second choice) 2/5 = 8/30
Thus, final probability is 8/30+8/30 = 8/15.

XXXXX

Re: finding the number of wanted outcomes using combinations:
Deal with each group separately, then multiply. To get exactly one sour, you want 1 sour (out of 2), and one sweet (out of 4). thus, there are 2C1 =2 ways of choosing sours, and 4C1=4 ways of choosing sweet. Multiply the two groups to find the number of ways of choosing your wanted outcomes.

For example, the same question with 4 sweet and 3 sours, where we need to choose 4 candies, prob of 2 candies sour:

Total number of outcomes: choose 4 out of 7, or 7!/3!4!

number of wanted outcomes - you basically want two sour, two sweet. deal with each group separately and multiply:
number of ways of choosing 2 sour out of 3: 3!/1!2!=3
number of ways of choosing 2 sweet out of 4: 4!/2!2! = 6
Thus, total number of wanted outcomes is 3*6=18.
(Basically, for each of the three ways of choosing the sours, there are 6 ways of choosing the sweets - hence 3*8).

Caveat2: You're not likely to see a question this complicated on the real GMAT.
Geva
Senior Instructor
Master GMAT
1-888-780-GMAT
https://www.mastergmat.com
Join the discussion

by diebeatsthegmat » Tue Sep 28, 2010 9:25 am
Jinglander wrote:I know how to solve this using a probability decision tree, but wanted to see if it could be solved with combinations.

Renee, has a bag of 6 candies, 4 are sweet 2 are sour. Jack picks two candies simultaneously and at random, what is the chance that exactly one candy he picked is sour.

I know the total combinations its just 6!/2!4!. But how do I find the number of outcomes that have exactly one sour. Clearly the numbers are small enough that I can just count the number of out comes but i want a formula that can be applied to larger numbers
if you this way, you will have to find the probability to find 1 sweet and 1 sour to get exact 1 sour candy.
the way to select 1 candy is 4!/3!1!=4 and select 1 sour candy is 2!/1!1!=2
total ways to select 1 sweet and 1 sour candy is 2*4=8
thus the answer is 8/15 ( 15 is the combination you found from 6!/2!4!

you can solve the problem by another different way
select 1 sweet candy and 1 sour candy or select 1 sour first and 1 sweet candy later
thus the probability will be 4/6*2/5=8/30 and you must multiple 2 because there are 2 times depending which kind of candy you selected, sour or sweet
the answer is also 8/15
Join the discussion